OpenStudy (anonymous):

Which of the following is the inverse of f(x) = 6/x? y = x/6 y = 6x y = 6 - x f(x0 is its own inverse.

OpenStudy (anonymous):

To find the inverse switch x and y: y = 6/x INVERSE x = 6/y yx = 6 y = 6/x It is it's own inverse